OpenStudy (anonymous):

HEEELPP!!!! The point (x,-1) is 13 units from the point (3,11). What are the possible values of x? Answer: -2,8

OpenStudy (anonymous):

\[(x-3)^2+(-1-11)^2=13^2\] solve for \(x\)

OpenStudy (anonymous):

\[(x-3)^2+144=169\] \[(x-3)^2=25\] \(x-3=5\) or \(x-3=-5\) etc

OpenStudy (anonymous):

how do you find x?

OpenStudy (anonymous):

add 3

OpenStudy (anonymous):

So it's not 3 to power of 2 then for (x-3)^2?